### v1

@@ 6 Conclusion
In conclusion, we proposed AMF-MedIT, an efficient Align-Modulation-Fusion framework for integrating medical image and tabular data, to address the cross-modal disparities during fusion and noisy medical tabular inputs. At its core, the AMF module leverages streamlined modulation and prior knowledge to effectively coordinate cross-modal disparities, providing an accurate and adaptable fusion method for various downstream conditions. Additionally, FT-Mamba enhances the performance of the tabular encoder, achieving an optimal balance between performance and data efficiency. Furthermore, our interpretability study on image-tabular contrastive learning reveals the impact of encoder selection on multimodal pretraining. Our approach enhances data efficiency and mitigates reliance on large annotated datasets, offering a novel solution to challenges in medical multimodal fusion. We hope these findings will inspire advancements in medical multimodal models and their applications in healthcare.
@@
Congjing Yu: Conceptualization, Data curation, Methodology, Validation, Writing - original draft, Writing - review & editing. Jing Ye: Conceptualization, Validation, Writing - review & editing. Yang Liu: Conceptualization, Validation, Writing - review & editing. Xiaodong Zhang: Conceptualization, Supervision. Zhiyong Zhang: Conceptualization, Funding acquisition, Supervision, Writing - review & editing.
or transparency, we require corresponding authors to provide co-author contributions to the manuscript using the relevant CRediT roles. The CRediT taxonomy includes 14 different roles describing each contributor’s specific contribution to the scholarly output. The roles are: Conceptualization; Data curation; Formal analysis; Funding acquisition; Investigation; Methodology; Project administration; Resources; Software; Supervision; Validation; Visualization; Roles/Writing - original draft; and Writing - review & editing. Note that not all roles may apply to every manuscript, and authors may have contributed through multiple roles.
@@ A.2 Multimodal Pretraining
The pretraining procedures and augmentations followed those described in (Hager et al., 2023), where image augmentations were applied with a probability of 0.95 and tabular features were corrupted at a rate of 0.3. The optimizer used was Adam with a weight decay of $1e^{-4}$. The learning rate was selected based on validation performance, with possible values of $(3e^{-3},3e^{-4},3e^{-4})$. The model was trained for 500 epochs using a cosine annealing scheduler with a warmup period of 10 epochs.
A.3 Supervised Training and Fune-tuning
Considering the severe class imbalance in the OL3I dataset, where class 0 contains 7,784 samples and class 1 only 355 samples (a ratio of approximately 20:1), the weight factor $\alpha$ in the weighted CE loss was set to 0.1 for class 0 and 2 for class 1. Additionally, a weighted sampler strategy was applied to the OL3I dataset, where class 0 in the training set was downsampled by a factor of 0.5, and class 1 was oversampled by a factor of 5.
To determine the embedding dimension $D_{tab}$ for tabular encoders, we evaluated multiple candidate dimensions during supervised training on the OL3I dataset. For MLP and ResNet, the values tested were $\{128,256,512,1024\}$, and for FT-Transformer and FT-Mamba, $\{32,64,96,128,256,384\}$ were considered. During the fine-tuning of the AMF module, $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 5.0 to ensure that the density and leakage losses have a comparable magnitude and regularization strength to the cross-entropy loss. To optimize the training hyperparameters, we tested three learning rates $\{1e^{-2},1e^{-3},1e^{-4}\}$ using a fixed seed of 2022 for all supervised and fine-tuning experiments. In the fine-tuning of the cross-attention module, we observed severe overfitting and thus tested weight decay values of $\{1e^{-1},1e^{-2},1e^{-3},1e^{-4}\}$ to regularize training. In all other cases, the weight decay factor was fixed to 0. The optimal values were selected based on the validation performance.
@@ A.4 Quantifying the Modality Confidence Ratio
In our experiments, we assessed the confidence levels of unimodal feature vectors, $\textbf{Confidence}(\boldsymbol{z}_{img})$ and $\textbf{Confidence}(\boldsymbol{z}_{tab})$, based on their respective unimodal performance. Specifically, after multimodal pretraining, we constructed unimodal models by utilizing the pretrained encoders and fine-tuned them on downstream tasks with a frozen strategy. The modality confidence ratio $R_{\text{conf}}$ was then calculated as the ratio of the image-to-tabular unimodal model metrics on the testing set. The unimodal performance and the corresponding modality confidence ratio $R_{\text{conf}}$ are presented in Table A.1.
Table A.1: Unimodal performance of pretrained encoders and corresponding modality confidence ratio.
Image Encoder Tabular Encoder OL3I AUC (%) OL3I DVM ACC (%) DVM
Image Tabular ${R_{conf}}$ Image Tabular ${R_{conf}}$
ResNet-50 MLP 74.99 77.41 0.97 66.11 79.65 0.83
@@
FT-Mamba1 72.92 70.98 1.03 - - -
  • 1
    Perfomance on incomplete tabular data.

### v2

@@ 6 Conclusion
In conclusion, we proposed AMF-MedIT for integrating medical image and tabular data. At its core, the AMF module leverages streamlined modulation and prior knowledge to coordinate cross-modal disparities, providing an accurate and adaptable fusion strategy across diverse downstream conditions. Additionally, FT-Mamba serves as an efficient yet powerful tabular encoder, achieving a favorable balance between accuracy and data efficiency. Our interpretability analyses further reveal that encoder selection critically shapes multimodal contrastive pretraining, highlighting the essential role of FT-Mamba in multimodal feature representations. Beyond architectural contributions, AMF-MedIT improves robustness against noise and reduces reliance on large annotated datasets, offering a practical and scalable solution to the realities of clinical multimodal learning. We hope this framework will inspire advancements in medical multimodal models and their applications in healthcare.
@@
Congjing Yu: Conceptualization, Data curation, Methodology, Validation, Writing - original draft, Writing - review & editing. Jing Ye: Validation, Writing - review & editing. Yang Liu: Validation, Writing - review & editing. Xiaodong Zhang: Supervision. Zhiyong Zhang: Funding acquisition, Supervision, Writing - review & editing.
@@ A.2 Multimodal Pretraining
The pretraining procedures and augmentations followed those described in [6], where image augmentations were applied with a probability of 0.95 and tabular features were corrupted at a rate of 0.3. The optimizer used was Adam with a weight decay of $1e^{-4}$. The learning rate was selected based on validation performance, with possible values of $\{3e^{-3},3e^{-4},3e^{-4}\}$. The model was trained for 500 epochs using a cosine annealing scheduler with a warmup period of 10 epochs.
A.3 Downstream training
Considering the severe class imbalance in the OL3I dataset, where class 0 contains 7,784 samples and class 1 only 355 samples, the weight factor $\alpha$ in the weighted CE loss was set to 0.1 for class 0 and 2 for class 1. Additionally, a weighted sampler strategy was applied to the OL3I dataset, where class 0 in the training set was downsampled by a factor of 0.5, and class 1 was oversampled by a factor of 5.
During the fine-tuning of the AMF module, $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 5.0 to ensure that the magnitude and leakage losses have a comparable magnitude and regularization strength to the cross-entropy loss.
To determine the embedding dimension $D_{tab}$ for tabular encoders, we evaluated multiple candidate dimensions during supervised training on the OL3I dataset. For MLP and ResNet, the values tested were $\{128,256,512,1024\}$, and for FT-Transformer and FT-Mamba, $\{32,64,96,128,256,384\}$ were considered. Based on hyperparameter search, the tabular embedding dimension $D_{tab}$ was set to 256 for MLP and ResNet, and 64 for FT-Transformer and FT-Mamba. The number of encoder blocks was fixed to 2 across all tabular models, following the configurations in [6, 25].
To optimize the training hyperparameters, we tested three learning rates $\{1e^{-2},1e^{-3},1e^{-4}\}$ using a fixed seed of 2022 for all supervised and fine-tuning experiments. In the fine-tuning of the cross-attention module, we observed severe overfitting and thus tested weight decay values of $\{1e^{-1},1e^{-2},1e^{-3},1e^{-4}\}$ to regularize training. In all other cases, the weight decay factor was fixed to 0. The optimal values were selected based on the validation performance.
@@ A.4 Quantifying the Modality Confidence Ratio
In our experiments, we assessed the confidence levels of unimodal feature vectors based on their respective unimodal performance. Specifically, after multimodal pretraining, we constructed unimodal models by utilizing the pretrained encoders and fine-tuned them on downstream tasks with a frozen strategy. The modality confidence ratio $r_{\text{conf}}$ was then calculated as the ratio of the image-to-tabular unimodal model metrics on the testing set. The unimodal performance and the corresponding modality confidence ratio $r_{\text{conf}}$ are presented in Table A.1 and A.2.
Table A.1: Unimodal encoder performance and corresponding modality confidence ratios with different tabular encoders.
Tabular Encoder OL3I DVM
Image AUC (%) Tabular AUC (%) $r_{\text{conf}}$ Image ACC (%) Tabular ACC (%) $r_{\text{conf}}$
MLP 74.99 77.41 0.97 66.11 79.65 0.83
@@
Table A.2: Unimodal performance of pretrained encoders and modality confidence ratios under noisy conditions.
Tabular Missing Rate OL3I AUC (%) Image $\sigma$ OL3I AUC (%)
Image Tabular $r_{\text{conf}}$ Image Tabular $r_{\text{conf}}$
0.25 72.92 76.14 0.96 0.10 69.37 76.21 0.91
0.50 72.92 66.93 1.09 0.15 64.16 76.21 0.96
0.75 72.92 64.07 1.14 0.25 56.89 76.21 0.89
A.5 Model interpretation
IG estimates the contribution of each input feature by integrating gradient values along a straight-line path from a baseline input to the target sample [31]. In our setting, the baseline input was defined as the mean of all training tabular samples. IG scores were computed for all training samples with respect to the disease class, and the final feature importance values were averaged to mitigate randomness. For consistent comparison, each feature’s importance was normalized as a percentage of the total importance within the model, highlighting its relative influence during inference. We also applied Grad-CAM [32] to generate class activation maps (CAMs) from the pretrained multimodal models. For a consistent comparison, all models used a concatenation-based fusion framework to avoid confounding effects from fusion modules.